OpenStudy (anonymous):

single variable calculus . help me friends find the maclaurine series for (i) f(x)=1/1−x

OpenStudy (klimenkov):

\(\frac1{1-x}=1+x+x^2+x^3+..., |x|<1\)

OpenStudy (anonymous):

solution?

OpenStudy (klimenkov):

Yes.

OpenStudy (anonymous):

can you explain how to get it?

OpenStudy (klimenkov):

It is the sum of the geometric sequence.

OpenStudy (anonymous):

oh.. okay2. how to find taylor series then?

OpenStudy (anonymous):

do you know?

OpenStudy (klimenkov):

\(f(x)=f(x_0)+f'(x_0)(x-x_0)+\frac{f''(x-x_0)}{2!}(x-x_0)^2+\ldots+\frac{f^{(n)}(x_0)}{n!}(x-x_0)^n+\ldots\)

OpenStudy (sirm3d):

\[\huge f(x)=\sum_{n=0}^{+\infty} \frac{ f^n(a) }{ n! }(x-a)^n\] compute the value of \[\huge f^{(n)}(a)\]

OpenStudy (sirm3d):

where f^0(a) is the value of the function, f^1(a) is the value of the first derivative. For the maclaurin series, use a = 0.

OpenStudy (anonymous):

@klimenkov : thank you :D @sirm3d : tHANK yOU :D